A project has the following cash flows:

Year Cash Flow
0 –$ 16,300
1 7,000
2 8,300
3 6,800

   

What is the NPV at a discount rate of zero percent? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

  

  NPV $   

  

What is the NPV at a discount rate of 12 percent? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  NPV $   

  

What is the NPV at a discount rate of 19 percent? (Negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  NPV $   

  

What is the NPV at a discount rate of 28 percent? (Negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  

  NPV $   
Answer #1

NPV =PV of inflows -Initial Outflow

The NPV can be computed using the excel formula=NPV()

NPV@0%

The discount rate =0%

The formula used =NPV(0%,B3:B5)+B2 we get NPV as $5,800

NPV@12%

The formula used =NPV(12%,B3:B5)+B2 we get NPV as $1,406.81

NPV@19%

The formula used =NPV(19%,B3:B5)+B2 we get NPV as -$521.25

NPV@28%

The formula used =NPV(28%,B3:B5)+B2 we get NPV as -$2,522.84

BS Book1 - Excel (Product Activation Failed) X File Home Insert Page Layout Formulas Data Review View Developer Tell me what